You are here:
 

TPCG News Room

Terrebonne Parish Consolidated Government

 

BP Press Release for June 7, 2010.

Tuesday June 08, 2010 08:05 am - 2960 Views

Terrebonne Parish Information 1)  Coast Guard representatives in Cocodrie reported:

  • Personnel are conducting waterside surveillance for offshore and shoreline impact of oil.
  • Total personnel in Cocodrie: 764.
  • Weather is expected to be poor to marginal for skimming operations until June 7.
  • The US Coast Guard reports the three standby TOP HATs were recovered from the seabed for modifications and will be available in case a problem is encountered with the TOP HAT that is in use.
  • Only one of the four vents on the TOP HAT is closed.
  • Motor Vessel (MN) ENTERPRISE has recovered 10,496 barrels of oil and flared 22 million cubic feet of gas during the last 24 hours.
  • Sub-sea dispersant injection continues at the Blow Out Preventer (BOP) at a rate of 10 gallons per minute.
  • Drill Rig 00111 conducting the 1 st relief well, remains at 12,956 feet below the drill floor. Drill Rig 0011 conducting the 2nd relief well, is at 8,650 feet below the drill floor. The Drill Rig 0011 test on the BOP failed a function test and is being brought to the surface for troubleshooting.
Boats/Vessels
  • Deployment of Skimmers: 1 positioned between Wine Island and Trinity Island, 1 between Wine Island and Timbalier Island, 3 skimming designated routes in Lake Barre’ and Terrebonne Bay.
  • 136 boats on site assisting with cleanup in Cocodrie.                 Assignments:                                 Surveillance:                     2                                 Decontamination:              3                                 Task Forces 1-3:              105                                 Recovery:                         26  136 Total
Boom
  • Boom crew is now focusing on Tier 3 and Tier 4.  Booming plans are nearly 60% complete.
  • Ongoing boom inventory in Cocodrie staging area.  Latest boom numbers:   1)  Hard Boom:                 a)  Deployed 6/5/10:    9895’                 b)  Total Deployed:      173,955’   2)  Sorbent Boom:                 a)  Deployed 6/5/10:    0*                 b)  Total Deployed:      83,900’* *No change due to inclement weather
  • Reported that the containment boom and sea boom on the Bay side of Raccoon Island looked good, but on the east side, the containment boom and sea boom was torn and floating up on shore.
  • Reported from the Gulf side of Raccoon Island that the containment boom and sea boom was torn off and washed onto the shore.
  • Assessment Team reported from Lake Raccourci that crews were removing large amounts of absorbent boom that had previously been deployed in the area.  They will redeploy boom after the used boom is removed.
  • Report of (5) boats working on boom repair at Wine Island.  It was noted that the large sea boom was being deployed around Wine Island and that USCG personnel are in all of the boom boats. 
Cleanup Progress
  • Continued beach reconnaissance and manual recovery of discovered oil.
2)  Oil impact reports:
  • Reported from the Gulf side of Little Pass that there was no visible oil on the surface, no visible sheen, and no visible oil on the rip current.
  • Reported from the western bay side of Little Pass that there was a slight sheen on the surface, approximately 100 yards wide and elongated on the rip current. No other visible signs of oil or tar balls were reported at these coordinates, other than a mousse on the rip current.
  • Reported from the eastern end of West Timbalier Island that a SCAT team was working that consisted of EPA contractors and (1) DEQ personnel working along the western end of West Timbalier Island. New micro-droplets of oil were observed and reported, approximately ¼”- 4/6” maximum in diameter, stranded along the beach face.
  • Assessment Team reported from Lake Raccourci that crews were removing large amounts of absorbent boom that had previously been deployed in the area.  They will redeploy boom after the used boom is removed. 
  • Report from 19th Canal, West Timbalier Island, east of the compressor station, that there was fresh pooled oil behind the rocks on the bay side. Oil staining was noted along the rocks and oil-coated vegetation was noted for 350 yards.
3)  Vessels of Opportunity Program:
  • Anyone interested in applying to becoming a Vessels of Opportunity (VOO) or crew member should call 281-366-5511
Training Opportunities for Deck Hands and Beach Clean-up Workers
Local training sessions have been scheduled for residents interested in becoming employed for the response to the Deepwater Horizon oil spill. Residents interested in working on a beach clean-up crew or as a deck hand on a vessel responding to the spill can attend this free training. Training organizers on site will provide information on how to seek employment in these fields. 
Each location will feature two training sessions per day with the morning training at 9:00 A.M. and the afternoon training beginning at 2:00 P.M.  The training location schedule is as follows:
The Larose VFW Hall (located at 2322 VFW Ave., just off West 23rd St. in Larose) will host training sessions on Tuesday, June 8, and Tuesday, June 15.
The BP Annex (located at 5703 Hwy. 56 in Chauvin) will host training sessions on Wednesday, June 9, and Wednesday, June 16.
The Grand Isle Community Center (located at 3811 Hwy. 1 in Grand Isle) will host training sessions on Saturday, June 12, and Saturday June 19.

Additional Information

Contact

Hours of Operation

  • Monday thru Friday
  • 8:00 a.m. to 4:30 p.m.

Mailing Address

  • P. O. Box 2768
  • Houma, LA 70361
 
Return To Top